Space War

Food For Droid Thought

by William Stein February 9, 2002

2

Contents

1 Droid Armies 7 2 Sundance Galaxy 15 3 Tai Stargate 21 4 Kelly Wong 29 5 Droid Space 35 4 CONTENTS

Preface

I wrote a first draft of a book called “Danger of Droids” when I was ten years old. The book you are looking at began as a project to rewrite and expand the original, however the two books diverge in the middle of chapter 3. It’s not quite done yet, though I know how it will end.

If you like this book, let me know.

I give anyone the right to read and distribute this book unchanged and for free in electronic form. However, the copyright remains mine.

William Stein http://modular.fas.harvard.edu 6 CONTENTS

Chapter 1

Droid Armies

Crisp cold air sweeps across a gray sky. A sturdy black building sits in the middle of a huge red and rocky area on a freezing barren planet. Numerous other identical menacing buildings are stationed across this dreadful landscape. Each contains something that only the recently appointed president, Columbo, and a few of his trusted advisers knows about.

* * *

Bill Stein recently celebrated his 25th birthday. He is about five feet ten inches tall, has black hair, and a medium build. At the moment, he is standing on a barren planet, near a dark precipice. He thinks, then looks at his watch, which, for a very brief moment, reads “2485, August 16, 1:25, 16 and 28/1000000 seconds.” Upon seeing the time, he pulls a gadget from his pocket and speaks ls into it. “This is Captain Bill Stein on Sector 4. Bring me back from this scout mission. It’s getting cold out here.” In fact, the cold can’t touch him in his heated space suit.

Far away, floating in space, the beaming office replies: “OK. I’ll beam you aboard the ship.” The beaming officer was quite unusual and harbored a deeply veiled past. In less than a second, Bill materializes aboard the ship. He removes his protective suit, exits the beaming room, and enters a hallway where he is met by a 6 foot 2 inch tall man with fluffy brown hair and sharp features. Together they step into an elevating room and are pulled through an antigravity tube. As their journey ends and the tube opens, Thom steps onto the bridge.

Thom is Bill’s best friend, and like Bill he is 25 years old. He stands just over six feet tall in front of his monitoring and communications array and in a few moments yells “Bill, look at the scanning display!”

“What exactly do you see?” Bill inquires.

“A blip,” Thom excitedly answers. This is Thom’s 6th mission.

“There’s something out there and I want to know what it is and where it is going. What is its present course?” Bill asks carefully.

“It’s on a direct course to intercept us. I may be able to bring up a visual of the blip using our ship’s telescopes”, Thom suggests.

“Try,” Bill agrees. Several barely-visible specs appear in the middle of a large black viewscreen. They wiggle a little, indicating rapid movement. “Thom, try to zoom in further and enhance.”

The view on the screen barely changes, as Thom studies his instruments. “Bill, those are sophisticated space ships, and they do not look friendly. Their weapons are fully charged.” Thom fixes a concerned eye on Bill, “I suggest we arm our lasers as a safety precaution.” The mood on the bridge becomes tense. Officers to the left and right of Bill try to appear to relax as they study their consoles while nervously glancing at the viewscreen.

“I agree.” Bill pauses, then turns to an officer nearer the view screen and adds, “Dennis, arm our weapons array.” Dennis is 24 years old and 5 feet 9 inches tall. If anyone really knows Bill, it is Dennis. Though Bill is captain of the ship, and Dennis is devastating in combat situations, without Dennis’s financial wizardry, the ship would never have been allowed just the right crew.

“Laser’s armed”, Dennis replies as his hands dance in the air amongst a forbiddingly colored hologram that hovers in front of his chair.

A flash of light appears on the viewscreen. A single thought leaves each officer’s mind as rapidly as it arrives: Will I be incinerated in a moment? The bridge shakes enough to knock Bill and Thom to their knees as a bolt of energy cuts into the ship’s launch pad!

“Fire lasers!” Bill yells. The air vibrates slightly as Dennis releases laser cannon fire at ten approaching ships. The viewscreen shows three fighter ships explode, as the shields on the other seven re-modulate in order to deflect Bill’s lasers.

“I suggest I try to establish communication with them”, suggests Thom. “Go ahead and try”, approves Bill. Thom tries, but has no luck. The

enemy ignores all hails. Frustrated and worried as the remaining seven ships speed closer, Bill orders “Ready the ship to jump to light speed.”

“All right” responds the lead cruise controller, a clever blonde haired young man named Chris, who sits to Bill’s left. The controller and three nearby officers move their hands carefully above a smooth panel admists rotating holographic images. The ship hums and vibrates as the hyperdrive engines are activated. Just as the ship begins to jump to light speed, enemy laser fire storms all around damaging the hyperdrive. “Captain, we’re not going to be able to jump to light speed for a while. I will try to assess the damage caused by those last blasts.”

“How long do you think it will take?” Bill demands.

“It will probably take a few hours,” replies the cruise controller slowly, keenly aware that he might not live more than a few more minutes.

“We’ll just have to fight them off” shouts Bill; “keep firing lasers Dennis, Thom, and Chris!” He mentions manuever “alpha bravo” to the cruise controller.

As the battle rages on, it rapidly becomes clear that Dennis is the best shooter. A total of six enemy ships fall by his hand. After a while, it seems that all of the enemy ships are gone, as the ship speeds seemingly unpursued across the deeply empty reaches of space.

Bill congrulates his crew: “Great shooting Dennis!” “Excellent work repairing our light-speed hyperdrive, Chris.” Then Bill thinks fleetingly about his lover Tasha, who was far away doing research near the Jash Stargate, and who he has not seen in six months. Was she safe?

A few tense hours pass, in which the crew carefully monitors for attack craft, and does their best to hide the weakened ship behind floating debri. Finally, Chris announces “The hyperdrive should now be fully operational.”

“Great!” replies Bill, and then adds solemly, “Let’s go home.”

“Just give the order, Bill”, Chris says excitedly.

“It’s given.” As the ship enters an open expanse of space, it hums and vibrates, and the hyperdrive safely jumps the ship to near light speed. For a moment every living thing on the ship is mildly disoriented, and the next moment the only change is that the view out the window is bizarrely distorted. Several hours pass, and then Chris slows the ship down, as most of the crew’s home world appears, as a dot, on the viewscreen.

“It is fabulous to see Earth again,” comments Thom, “When we were attacked we fought well; I don’t think we can easily be captured.”

The ship rapidly approaches Earth, shooting through space at sublight speeds that defy comprehension. As the planet grows to encompass most of the viewscreen, Bill thinks for a moment about his parents below, and wishes that his lover Tasha were also below instead of on a research mission at Dinesta. As the ship breaks through the automosphere and glides towards Earth, a landing area appears, imprisoned on three sides by water. Thom notices that the fourth side is populated with an organized large number of people and asks, “What is going on down there?”

“I don’t have the slightest idea”, Bill answers. “Chris, slow our landing. Thom, run a scan on the people who appeared armed with laser cannons.”

Moments later, Thom yells in surprise: “Those ’guys’ are droids!”

“They must be somehow connected with the droids that I discovered in the sheds on that planet’s surface,” realizes Bill.

The ship jerks, knocking Bill and Thom temporarily off balance. “What just caught us?” asks Dennis.

“Some kind of tractor beam”, responds Thom.

Before Bill can say “ready lasers”, the ship is locked onto the landing pad, and the weapons array is covered by special mirrors, that are usually placed there to prevent accidents, but in this case make the ships weapons all but useless.

“Let’s try to break free”, Bill cries. Chris tries, but the ship is held fast. “Let’s try to shoot the droids”, suggests Dennis. “No – with the safety deflectors over our lasers, we’ll only harm ourselves,”

warns Bill.

Red lights flash as the ships goes into battle alert. Droids break through an entrance and begin boarding the huge ship. Some are ripped apart by laser fire from crew members, but soon the droids produce shields which effectively protect them.

Feeling betrayed by the security of their home world, eventually, Bill, Thom, Dennis, and Chris all find themselves being forceably escorted off their ship. The rest of the crew is detained onboard by ever patient and watchful droids.

“Why are you here?” inquires Bill as he and the others are moved very quickly away from the ship in a carefully sealed rectangular black vehicle. The droids take no notice. The vehicle connects solidly with a featureless grey building, and unloads its human cargo. The human-looking expressionless droids lock our space travelers up in a medium size featureless cell. The cell walls are smooth and grey, there is a bench against a wall, and a small glowing force-field door, through which nothing can be seen.

“Thom, just moments ago you said you didn’t think we could be beaten or even captured”, says Dennis in frustration.

Thom stoically replies, “we are not dead yet.”

“We are in a big mess,” Chris notes. “We have to escape and get help from one of the allied planets.”

“We had better do so,” agrees Bill, as a droid approached the force-field door of their cell.

“Be silent”, says the guard in a monotone voice from behind the door. The sound of the guard’s footsteps suggest that he walks several yards directly away from the door.

“I have a great idea!” announces Bill very quitely.

“What is it?” whispers Tom.

“Well, we could probably make it past several of the guards...”, he lowers his voice even more, “if we could re-program one of the other droids. We’ll make a lot of noise to draw the guard’s attention. It will come in here to shut us up, and then Dennis will take him out. Dennis, I trust you were able to smuggle off of our ship something that could be used as a weapon?”

“Of course.” Dennis confirms, with a twinkle in his eye as he taps an almost imperceptible bulge in the corner of his pantleg.

At Bill’s signal, Chris, Thom, and Dennis begin talking, then yelling loudly, and even appear to be fighting each other at one point. The guard droid, who had warned them before, appears stoically at the door and says “Be silent.” A torrent of loud curses echo back from the small prison. The guard, devoid of fear, follows its programming; it shuts down the force door. It enters the room, then the force door is re-activated by another guard.

Dennis springs from behind at the guard, brandishing nothing but a thin diamond wrench, and makes a quick incision from behind. The guard stops in midstep and falls to the ground. Everyone is silent; they must work quickly.

Thom whispers, “Assuming Dennis hasn’t damaged this droid too much, I’ll try to re-program it so that it will attack the other droids when they least expect it. It’s too bad we don’t have more time to study these droids, so we could perhaps learn their purpose.” Because the droid looks so human, it is disconcerting to see Thom open various panels and beginning hacking through, then reprogramming the mind of the droid, using a holographic interface that he is able to produce.

“Hurry up” Bill whispers in a concerned voice, “I hear footsteps. The other droids will soon become suspicious.”

The force door opens, revealing two droids, just as Thom’s droid stands tall and announces in a convincing monotone, spoken in the droid’s own encrypted language, that the prisoners are under control, because he has just administered a sedative to each of them. Thom feigns to look tired, and Bill, Dennis, and Chris follow suit. As the droids turn and begin to leave, Thom’s guard droid fires poinpoint red lasers from each thumb, which land squarely in the backs of each of the other two droids. They stop moving, then fall to the ground.

“Amazing work” congrulates Bill. The four jump through the force door and into a hallway of Building 7, a building they all got to know well during their early years of training. “We’re near what used to be a weapons storage room; let’s check to see if it still is.” Bill leads the way, followed by Chris, Dennis, and Thom and his droid.

They quickly find the storage room, while appearing to avoid detection. The storage room is locked by a black force shield. Because he is on the outside of the force shield, Thom has little difficulty breaking through it using a “backdoor” that he installed while in training here a few years ago. The force shield dissipates, revealing stocks of weapons lined up against the walls and a suspicious and heavily armed droid in the center of the room, who immediately begins to fire as our four fighters jump away from the door. Last blasts darken and rip apart the side of the hallway opposite the door.

“Surrender or die” the attacking droid threatens as he emerges from the weapons room. Just as our fighters are about to run, Thom’s droid makes short order of the attacker with a quick small red laser blast from his right thumb. Moments later, our fighters are armed to the teeth. They charge down a corridor, with hopes of amassing enough power to retake their ship and escape.

“Watch out behind you, Bill! Duck and fire!” screams Thom. A laser blast barely misses Bill, resulting in a mild burn on his arm, as he fires back hitting nothing. Shots from Thom hit the attacking droid, causing it to explode.

Eventually, a systematics search leads the crew to the droid’s command and control center, which is located in a small tower mounted on the top of Building 7. With some difficulty, they manage to take the tower. Though Thom doesn’t understand the encrypted wireless protocol used by the droids to communicate, he can send messages using it via his loyal droid, who has managed to escape harm so far. Thom sends out a message using the droid’s encrypted wireless frequency telling all droids that they “must go to landing strip 14 immediately!” Even his own droid finds itself compelled to obey.

As Thom, Dennis, Chris, and Bill reboard their ship, they find their crew mostly unharmed and in the process of mounting a rescue effort, and have successfully disabled the ship landing constraints. When they reach the bridge, Chris says, “The controls appear to be intact. I think I can get us out of here. Just give the order, Bill.”

“It’s given,” says Bill, and the ship shoots quickly into the sky, as their view of the base disappears behind clouds, and the whole planet appears to them as little more than a dot. As some droid-controlled ships begin to fire on them, they engage their hyperdrive and are gone.

“That was very close,” comments Thom. “Let’s drop out of hyperwarp, then approach the nearest allied planet and explain to them what has happened to Earth.”

“That’s just what I had in mind,” agrees Bill. Chris slows the ship down, as Thom determines their exact location in the astromap database.

After a few tense minutes, Thom announces “The closest planet is Drizzt. Unfortunately Bill, my scans suggest that Drizzt is crawling with droids, just like Earth was.”

Just then a squadron of small agile attack ships is detected by their sensors. The approaching ships are barely visible by the light of Drizzt’s sun. They are mostly black with sharp streaks of silver running from their front to back. Each ship carries several huge red laser canons mounted on top and bottom. These ships have one purpose, to destroy.

Dennis yells, “They’re firing at us!! Fortunately our shields are fully charged, and have held so far.”

“Return fire!” Bill commands.

Two attacking ships explode moments after Dennis targets the attack ships and causes his weapons to release a few blasts.

“What happened to the lights? Ouch!” yells members of the crew. The remaining attack ships had simultaneously released a huge torrent of lasers and struck a direct hit on the ship’s primary power reactor.

Thom says in frustration, “Our weapons likely won’t be operational for the next hour.”

“No! I won’t give up so easily,” responds Bill, “Dennis, switch to secondary power. Are any of our weapons now at least partially operational?”

“We still have two of our laser cannons”, replies Dennis, as he prepares to fire one of them. “Bill, you may want to fire the other cannon.”

A small explosion rocks the ship. “Damn!”, yells Dennis, “I had an enemy ship targeted and just as I fired my cannon failed!”

“Use mine”, Bill replies. The lights re-appear, but are dim. “Thom and Chris, how close are we to having enough power to engage our hyperdrive?”

“Probably a few more minutes”, responds Chris as he puzzles over a holographic image that floats before him.

As three enemy ships close in for a final assault, the ship activates its hyperdrive, and, as far as the three ships are concerned, is gone. The droids who control the remaining attack ships fly back to planet Drizzt.

Sundance Galaxy

Now that we survived the fast-paced present-tense first chapter, the author has elected to switch to the past tense.

After the crew got some rest on board, and attended to their wounds, they partially repaired the ship. They then slowed from hyperspeed as they entered the Sundance galaxy.

Several of the ship’s top officers gathered together in the strategy room, a cool expansive high-ceiling room constructed in the ancient European style, with genuine stone floors and walls, and detailed woodwork decorating the walls. The holographic computer terminals, which looked out of place mounted on the walls, gave off a noticeable glow in the dim light.

When Dennis, Thom, Chris, and a few other officers had arrived, Bill spoke up: “We have lost all contact with the people or military forces of Earth. We don’t know what happened to our people, if they are alive, or if they are being held captive. The worlds neighboring Earth appear to have met a similar fate. One thing is for certain—we must retake the Earth!”

“We could try to convince some of the powers based at the planets in Sundance to help us,” Dennis suggested. “It would be in their interest, because their worlds might be attacked by the droids next.”

Thom added, “I made some preliminary scans, which suggest that perhaps the droids have not yet infiltrated Sundance.”

“Repairs on our ship are proceeding well,” noted Chris.

Bill turned to the chief engineer, a short and clever man, and said “Good work on the repairs, Nathaniel.” He paused, then said “Dennis, Thom, and Chris, meet me on the bridge.” The meeting ended.

“Thom, have you been able to establish communications with any of the leaders of nearby planets?” inquired Bill.

Thom responded affirmatively: “Davitt, leader of an influential nearby planet named Orbis, has agreed to speak with us in one hour. In the past, Orbisians have been sympathetic to human problems; moreover, they have a historic distaste for droids.” The hour passed uneventfully as repairs on the ship continued, and the crew did its best to review the place of Planet Orbis in recent Sundance Galaxy history. Finally, Davitt appeared on the viewscreen. The ship and Davitt communicated via a high-bandwidth encrypted link, which kept Davitt’s location quite vague.

What they saw on the screen did not look anything like a human. Daviss was quite a bit larger than a human, and wore a black and mirrored suit that allowed him to hover gently above the rough floor. His head hid behind a helmet that betrayed nothing of his facial expression while keeping him extremely comfortable in almost any atmosphere. Davitt began, his words automatically being translated to Earthlish, “My secretary has informed me that an army of droids have attacked Earth. This is consistent with reports we’ve received from our intelligence officers.”

“Orbis could be next,” responded Bill, very much to the point.

“A droid army once attacked us, and almost managed to imprison and slaughter much of our population. Fortunately, most of the planets of Sundance were united, and we were strong enough to repulse them. They must have retreated towards the Milky Way. At the time, we were not strong enough to follow them and destroy them completely.” He paused for a moment, “I have to wear this suit at all times because I was greviously injured when a spy droid attempted to assassinate me. These machines are terrible; they’re a virus.”

This recent-history lesson from Davitt served to chill Bill’s crew, as their thoughts turned once again towards their families and questionable futures.

“I will send forces to rendevous with you,” said Davitt to Bill, then he terminated the communication. Davitt, the absolute leader of Orbis, contacted his favorite and most trusted general, “Jerrisons, the time has come for us to launch a counterstrike against the droid armies. I have strong evidence that the droids have taken Earth; if so, they once again pose a supreme threat to us. We must strike now. Lead several rapid-action squadrons of our best fighters and rendevous with the Earth ship that you’ve no doubt detected by now.”

After a few hours of preparation, four elite squadrons take flight, breaking through Orbis’s gravitational pull, and begin flying towards the Earth ship. Half way into their journey, and without warning, the two lead ships are attacked!! “Watch out fighter K713, there is an unidentified fighter approaching you from behind!”, warns Masas, one of the fighter pilots, “I don’t detect any life signs on board, so maybe it’s a droid scout ship. Permission to destroy it?”

“Granted”, replies the occupant of K713, as he opens fire. Div-div-div-div-div-boooom!!! yell the lasers of K713, as the attacking ship is fragmented into millions of pieces. “Watch out Masas!” yells the occupant of K713. But he is too late. Masas’s ship is ripped apart under the assault of scorching merciless laser blasts, which issue from a second droid scout ship. The second ship then fires on K713, which spins and jerks as a laser blast glances off one of its stabilizer arrays. “I’ve been hit,” shouts the occupant of K713, as he sends out a distress signal. The droid closes in for termination.

∗  ∗  ∗

Back on the ship, Bill was concerned. “I wonder why it is taking Davitt’s forces so long to reach us?”

“I just detected a distress signal!” replied Thom, “It appears to be coming from one of Davitt’s ships.”

“There’s been recent laser fire,” chimes in Dennis, “it looks like one of two point fighters have been destroyed, and the other, K713, is disabled. I recommend we beam the pilot of K713 out immediately.”

“Do it!”, shouted Bill; the ordered propogated to the mysterious beaming officer, and, as a droid laser blast ripped apart K713, the pilot materialized in the beaming room, pleasantly surprised to be alive.

“Dennis, try to destroy the droid fighter, before it attacks any more of Davitt’s forces,” commanded Bill.

“Yes, sir.” Dennis launched an expensive unmanned craft, which, within ten minutes, efficiently sought and destroyed the loan droid scout in a kamakaze attack that resulted in its own destruction.

Within 20 minutes, Davitt’s four squadrons came into view, illuminated by the distant warm glow of Sundance’s trio of suns. The huge recently-saved occupant of K713 stepped slowly onto the bridge. “I am Jerrisons. I owe you my life.”

“It is honor to meet you,” shouted Bill. Jerrisons’s imposing stature and thick black battle armor made it difficult to avoid speaking loudly to him. “Let’s discuss our plans to regain Earth.” The two stepped off the bridge, followed by Thom and Dennis, and began to discuss the possibility of using the ship’s hyperfield to transfer the Orbis fighter squadrons near Earth. The Orbis military had not successfully deployed hyperdrives, because they chose to keep focused on their home world and neighboring planets, instead of scattering their attention throughout the Universe. Thus they were able to temporarily stem off the droid invaders but were not able to arrest their retreat. Earth, on the other hand, had the exact opposite focus, and had quickly fallen to the droids, as ship’s such as Bill’s would sadly discover as they returned home from fantastic voyages and were imprisoned by the droids.

∗  ∗  ∗

“Position your fighters at the following coordinates ...” commanded Colum, the supreme droid master whose thin stature towered over 9 feet high. He was the creator of the droid armies. He looked somewhat human, despite his height, and that his smooth hard skin was perfectly black.

“Yes master,” answered the droid in an encrypted transmission that is best described as an obedient monotone by those who frequently listen to such transmissions. Colum, who was born several hundred years ago in a distant ancient world, wanted to be present at this supreme moment of victory. Finally, Earth and any planet near enough to matter, would be completely under the control of the self-replicating droid army that he had built over a hundred years ago to help him liberate a new place for his people. All that remained was to take care of the last of Earth’s remaining annoying Space Voyagers, the Stein Ship. Being here was one of the greatest risks Colum had ever taken, but he needed to be certain of success.

Every so often a stroke of cosmicly bad luck visits a very powerful being. It would soon visit Colum.

Colum knew that Bill’s ship must pass near here in order to return to Earth, because this was the only safe stargate into the Milky Way. (Bill’s ship could activate its hyperdrive anywhere, but had to exit at a stargate.) Colum knew that Bill must come soon, as he had received a signal from two of his droid scouts ship. I will finally have them, Colum thought to himself as he rubbed his perfectly black four-fingered hands together.

Bus, Colum’s officer and friend for hundreds of years, sent a signal: “Colum, there’s an asteroid storm heading straight for you!!!!”

“What??” Colum’s ship was impacted repeatedly by a flurry of huge and pebble-sized ancient rocks which, billions of years ago, had started on just the trajectory that would eventually lead to a surprise attack on Colum’s ship.

Colum’s ship was ripped apart as Colum let loose a last angry scream of supreme anger and frustration.

Bus was unharmed, and realized this was his chance to take command of the droid armies, which is something he had longed to do for many many years. Bus brought his eight finger tips together and smiled. “What amazing luck.” There was just this small matter of the renegade human ship to take care of.

Bus opened a communication channel to the nearby droids, who had stopped forming their trap when Colum’s ship was destroyed: “Colum was terminated by an asteroid storm. Droid tactical squads, resume moving into ambush formation.”

“Yes, yes, yes, ...”, responded the countless droid fighters, in a lifeless monotone.

Bus extended his communication to all droids; his messages would quickly propogate throughout the galaxy: “Droids, in the last moment before Colum terminated, he made me supreme leader of the droid armies.” The droids believed him, though what he said was a big fat lusty lie.

* * *

“We should exit the stargate very soon”, announced Chris, “the hyperdrive is functioning perfectly, and we can already begin to see outside of the stargate.”

Bill turned to Thom, “Jerrisons has told me some of what he knows about how the droid army operates. Using only a skeletal force, they enter a galaxy and conquer a distant sparsely populated planet or moon, then spend a few years creating a droid army manufacturing facility that turns the raw resources of their base into droid fighters. They then have a constant supply of reinforcements. The Orbisians were lucky to detect this operation in Sundance Galaxy and head it off before it was too late.”

Thom responed, “We however, did not. The droids must have established such a base somewhere in our galaxy. I’ll immediately set a crew to search for where the droids have hidden their operations base. Perhaps they have become lazy in hiding it, because they seem to have all but taken complete control of our galaxy.”

As the star gate drew closer, Bill said, “Jerrisons, make sure your squadrons are ready for battle. It’s likely that we will encounter hostiles near the stargate. All fighter squadron, red alert!” On the viewscreen Bill could see a huge field of slowly-moving asteroids on the other side of the stargate, along with a mysterious shiny silver planet called Josh. Though many scout ships and probes had landed on the planet, it was believed that none had ever returned nor even sent back data.

The ship shook and vibrated firmly as Chris gently guided it out of hyperspeed and through the stargate. Bill’s forces entered the Milky Way.

Bus sent an encrypted signal to his forces, “Droids, wait until the humans are half way through the asteroid belt before attacking. They will fall into our red hot hands!”

“We are ready,” the droids replied obediently.

Tai Stargate

“Thom, what have you found?” asked Bill, “Have our scans picked up any droid ships?”

“Oh yes,” responded Thom in a deep voice. “They’re everywhere, though I can’t get a visual on any of them. They are hiding in the asteroid field.”

Bill replied, “So, they mean to ambush us. I wonder if they expected us to return with several squadrons of rapid-action Orbis fighters?”

From the bridge of Bill’s ship, Jerrisons was in constant communication with his squadrons, who had by now separated from the ship and were carefully flying in tight formation very nearby. Dennis had errected a powerful force shield near the ship, which protected both it and Jerrisons’s fighters, but which would reflect any weapons Dennis fired back. They were relatively safe from laser fire as long as the shield held out; unfortunately, the effect on their weapons was like the effect of lodging a piece of lead in the chamber of a gun.

Bill tried to break the tension as he inquired, “Thom, now that we are a back in the Milky Way, has your team had any luck searching for a droid manufacturing station.”

“Let me check.” There was a pause. “Maybe”, responded Thom, “there’s a planet called Roxy near the Tai Stargate. My team’s searches of Galaxynet suggests Roxy is the most likely home of a droid manufacture point. I think Chris will agree that we could re-enter the stargate and make a jump to Tai.”

“Let’s wait,” responded Bill. “Dennis, raise shields around both us and Jerrisons’s forces. Thom, any sign of the exact location of our ’welcome party’?”

Bus sent the signal: “Attack!!” Laser bolts exploded from several directions.

“Our shields are holding,” reported Dennis.

“I could hyperwarp us towards Tai Stargate in only a few minutes,” added Chris.

“Droid ship 6K7, kamakazi into the Stein Ship,” commanded Bus, as he began to test his enemy’s defenses.

“Yes,” replied the droid occupant of 6K7 obediently as his ship dived towards Bill’s ship. Bus suspected correctly that Bill’s ship’s shields would deflect energy bolts, but not a fighter ship. Expendible pilots and a selfgenerating army provided a formidible advantage against even the most cunning foe.

One of Jerrisons’s fighters, the 7th Falcon, broke from formation, and shot out from the shield, intercepted 6K7, and destroyed it, then took heavy fire as he retreated. The integrity of 7th Falcon began to fail, as it re-entered Bill’s shields. The pilot ejected, then was beamed aboard Bill’s ship, just before the 7th Falcon imploded.

“Bill, I think the droid leader’s is part of this ambush,” shouted Thom. “And, I bet I’ve located it. What a brilliant stroke of luck. ”

Bill responded, “Or a brilliant trap? I think not. Chris maneuver our ship towards the coordinates Thom gives you. Dennis, prepare to lower the shields and launch our photon disintegrators.”

The mood on the ship was tense. Combat always made Bill uneasy. His crew’s primary mission was exploration, not combat. They would much rather understand the universe than save it.

“Fire!” yelled Bill, as Dennis unleashed a torrent of photon disintegrators at a nearby asteroid. As the asteroid turned to dust, Bill’s crew saw a terrifying black and flourescent orange ship appear before them. It dwarfed Jerrisons’s rapid fighter ships, but was in turn dwarfed by the Stein Ship. Then it vanished, as the Stein Ship raised its shields.

Dennis summarized the situation, “Bill, that ship’s activated some sort of alien cloaking device. Even if we lower our shields, I probably won’t be able to lock any of our weapons onto it. If Jerrisons’s fighters start shooting wildly for it, they’ll be picked off by the other droids.” Dennis stood up from his chair to say something, but then changed his mind.

“Dennis, what is it?” inquired Thom, concerned.

“Oh, nothing,” answered Dennis, as if he had just awoken from a deep sleep. Dennis escaped Thom’s scrutiny as he began manipulating the rainbow of holograms that danced before him.

Chris interrupted, “Several droid ships are launching kamikazi attacks on us.”

“Jerrison, reattach your fighters to our ship. Chris, get us out of here! Let’s go to Tai Stargate as soon as possible.”

The droid ships caused some damage as Chris piloted the ship towards the stargate. Two of Jerrisons’s fighters and their pilots were destroyed defending the strategic retreat. “I’m ready to activate hyperdrive, Bill,” said Chris, “just give that sweet order.”

“It’s has just been given,” said Bill as the ship jerked then fled through the stargate.

Bus’s ship frustratedly decloaked. “That was good thinking,” praised one of the best-constructed droids. This droid believed it also could think.

“It was not good enough, I’m afraid,” responded Bus in what might be interpreted as a deep alien voice. “I should have never considered your suggestions during the battle. Perhaps droids can’t be programmed to really want victory.”

“What is the problem?” inquired the droid.

“These remaining humans are probably going to destroy our droid manufacturing sheds. If you would have monitored their Galaxynet traffic you wouldn’t have to ask. I only hope we can stop them. It is a shame that Columbo’s ship, which was destroyed, was equipped with hyperdrive technology, but ours is not.”

* * *

Chris broke the silence on board the bridge, “We will arrive at Tai Stargate very soon.” The ship slowed as it approached the stargate exit. Every eye on the bridge was fixed on the viewscreen as the ship exited the stargate and Planet Roxy came into view. A redish planet appeared on the view screen. Chris kept the ship far from the planet, in hopes that they cold avoid being detected. Planet Roxy was not pretty; she had a deadly atmosphere and her surface was as cold as the glacier capped polar regions of Earth.

Bill wasted no time, “This planet has ideal resources for a droid manufacture plant, and our sensors seem to have found one. Thom, Dennis, and I will beam down and try to sabotage the manufacturing facility.” Bill turned and said, “Dennis and Thom, meet me in the transporter room after you put your life-support suits on.”

Bill left the bridge, and passed Nathaniel on the way. Recalling Nathaniel’s facility in engineering destruction, he said “Nathaniel, you’re going down to Roxy with us.” Bill and Nathaniel both entered the antigravity tube, dawned life-support suits and arrive at the transporter room. “Is everybody’s suit operational?” asks Bill, who added, “Enter the life-support test chamber.” Thom, Dennis, Nathaniel, and Bill entered a special room that simulated conditions on the planet. It was imperative that if one of their sophisticated bulky suits was not defective or not configured properly that the mistake be discovered on the ship rather than under enemy fire on a hostile planet. Everbody’s suits checked out, and each person stepped, somewhat clumsily, to the transporter pads.

Thom said, “We should now exchange maximally secure public encryption keys so that we will be able to communicate during the mission with minimal chance of being understood.” He then clumsily turned his helmet enclosed head towards the beaming officer and asked, “Are you ready to beam us down to the planet’s surface?”

“I’ll be ready soon,” replied the beaming officer as he manipulated the multi-colored holograms that floated above his control panel, “I don’t want to beam any of you into something solid, you know.”

Bill, Thom, Dennis, and Nathaniel prepared themselves mentally for their critical mission as the beaming officer made a few last minute calculations. Finally, all was ready. The beaming officer slowly moved has hand through the hologram which caused one spatial dimension of Bill’s team to change ever so slightly. They vanished from view, just as the square leaving Flatland vanished. With that spatial dimension adjusted, the bizarre curvature allowed their usual position to be almost immediately altered. It was good that the movement occured so quickly, so that their bodies didn’t begin to break down. Transporting this way was OK for relatively small distances, but a voyage across a galaxy would destroy the voyager.

The team soon appeared on Roxy’s inhospitable surface. “This place is creepy,” said Nathaniel as he surveyed the rocky red landscape of the windswept freezing planet. The harsh expanse of scenery was interupted by a large mostly featureless black buildings.

“Abe, can your hear me?” inquired Bill, sending a signal back to his ship. Abe Burns, who was in charge of the ship in Bill’s absense, replied “Yes,

your signal is very strong. We’ll try to keep the transporter system locked on, and the ship ready to jump through Tai Stargate.” Abe, who stood just under six feet tall, was the same age as Dennis, and had studied with him many years ago on Earth. A concerned look swept across Abe’s face. He was more comfortable and happier while engaged in his research than when commanding a space exploration ship. Nonetheless, Abe was a superb and brilliant officer and Bill felt he could rely on him.

Back on the planet’s surface, the landing party planned their next move. Thom reported, “My sensors suggest that the entire planet can potentially activate a shield if danger is suspected. Thus if we move the ship near the planet and launch a volley of weapons, they will probably have no effect. Moreover, we’d be attacked by droid defense ships, of which there must be many as this is where they are manufactured.”

Nathaniel inquired, “Why don’t we simply beam a collection of timeactived megabombs onto the surface of the planet?”

Dennis responded “Extensive research has shown that they would probably explode the moment their other spatial dimension is modified, and our ship would be destroyed. Our equipment and bodies are able to stand a slight bit of spatial distortion but it’s far too risky to ask the same of a megabomb. I’d love such a weapon, but alas.”

Thom then suggested, “Nathaniel, if we can get you safely into one of these buildings, do you think you could cause it to self destruct?”

Nathaniel responded, “Sure. Not only that, but judging from the way our scans suggest these buildings are connected, I might even be able to start a carefully timed chain reaction, which would result in all of the buildings exploding almost simultaneously. That might even rip this god-forsaken planet apart!” This last sentence was pronounced with an almost disturbing sense of glee.

Dennis, Thom, and Bill were stunned. Bill asked, “Do you think you can really do that, Nathaniel?”

“Yes, I bet I can,” affirmed Nathaniel, “make sure our ship is ready to get us off this planet quickly.”

Abe’s studied some sensors on the bridge. He didn’t like what he saw. “Bill, you might be in trouble. The planet’s shield is activating.”

“I’m surprised, but a few droids are headed our way. Somehow they must have detected us. Try to keep our communications channel open,” responded Bill. Unfortunately, Abe’s reply could not be heard because of Planet Roxy’s shield. “Take cover behind this ledge, which leads to that building!” commanded Bill. “We’ll hold off these droids as we approach the building, then try to get inside so that Nathaniel can work his magic. Thom, keep trying to contact our ship. Dennis, take out as many droids as you can.”

Bolts of energy flared by Dennis’s head, as he peared through his handheld shield over the ledge, took aim, lowered the shield for a moment, and fired. A droid exploded in flames just as his shield re-appeared in time to protect him from another volley. “That was a little too close,” whispered Dennis as he became more cautious in his defense of the group and pulled from his pack his most advanced weaponry systems.

Soon Nathaniel, Bill, and Thom arrived at an entrance to the large black building, as Dennis continued to set up various high-tech defenses to hold of the droids. The entrance to the building was guarded by a black force shield, which was similar to but much stronger than the one that had imprisoned them on earth long ago. Thom tried directly to open the door via the control panel but was unsuccessful.

“We’re heavily armed. Let’s just blast the door away,” suggested Nathaniel, “Thom, stand back.” Nathaniel blasted away at the door several times, then Bill and Thom both joined in for a moment. The dark entrance didn’t appear to change, but Thom informed them that its strength had diminished approximately 10 percent.

Dennis noticed the situation and yelled “Stand way back and raise your body shields!” For a moment he turned from systematically picking of the droids, who were by now mostly pinned down in some cracks in the ground, adjusted his angry black and silver rifle, and fired one extended red bolt of energy which sailed slowly through the scorched atmosphere and injected itself into the portal. Dennis turned back to continue laying down defenses as the portal disolved. “I’ve got you guys covered out here. You can enter the building. I’ll let you know if the situation changes or if the droids send serious reforcements.”

“Let’s move in,” agreed Bill, as he led Thom and Nathaniel into the mysterious building. As the three passed through the door they entered a very dark corridor whose lights had all broken because the door had been ripped apart so badly that much of the atmosphere in the building had violently escaped. Fortunately their space suits were euipped with various lights that their face shields enhanced. They turned left and proceeded down the corridor. After walking for several minutes, worried at every moment that they’ll be attacked, Thom says: “This is strange. My sensor suggests faint signs of life nearby. Perhaps we are somehow throwing off the readings, or there is a native life form on this planet that thrives in this building.” The corridor turns right and they continue to follow it.

As they pass two heavy black force doors, Nathaniel boasts quietly, “After I get through with the control room we’ll be able to catch a glance in there on our way out.” Finally, ahead on the right, they spy what must be the control room. Four tense droids are just visible inside, and there is no door.

Thom says to Bill, “Droids don’t seem very intelligent. Perhaps if we make a noise they’ll all come out and we can shoot them. They’re expecting us anyways.” As Thom speaks, his sounds are completely muffled by his helmet. The sounds are received by a small microphone, encrypted, and broadcast via a weak signal that is usually lost to an eavesdropper in the ambient radio waves.

Dennis’s signal is a little stronger—“Bill, this is Dennis. The situation out here is becoming critical! At first the droids were easy to hold off, but they seem to be quickly adapting. I’ve only able to maintain our defense this long because of our superior weapons technology. You’re team better hurry up. Also, I’ve not been able to re-establish communications with Abe.”

Bill had little time to consider other possibilities and decided to follow Thom’s plan. They’re shots exploded towards the control room. As the smoked died down, there was no sign that the droids had relinquished their position in the control room, which deeply puzzled Thom. Suddenly a searing laser bolt ripped through the air from behind, and burned Nathaniel’s right arm. Thom and Bill returned fire as the approaching Droid tossed itself through the air shooting two laser pistols, an automated version of the Hong Kong action superstar Chow Yun Fat. A curse escaped Thom’s lips as he dodged to the ground. Time slowed as laser fire exploded everwhere. At last the droid’s shocking technique succumbed to Thom and Bill’s superior weapons. It shut off.

A similar scene ensued as the three desperately fought for the control room. As droids jumped through the air firing, the droids were eventually merciless ripped apart by Bill and Thom’s scatter guns. Nathaniel swore the droids fought for their “life”. “Something is wrong, very wrong,” he commented, as he began his work on the master control center. Just as promised, even with his right arm badly burned, he was able to shut down power to the droids in the building. He worked further on engineering a chain reaction that would hopefully destroy the planet while also allowing them to re-establish communication with Abe and beam to safety. After many very long tense minutes, Nathaniel at last announced: “I’m done. Once I set the reaction in motion there is no turning back. Probably this will allow us to contact Abe and get out of here. Just give the order Bill.”

A worried look hung to Bill’s face as he reassuringly spoke his trademark: “It’s given.”

Nathaniel moved his hands innocently, there was a warm buzz, and It began. “This planet will be history within 15 minutes,” said Nathaniel solemly, “let’s hope that we aren’t also.”

Just then, a pale but strong chinese woman stumbled out of the hallway and into the control room. She coughed and spasmed as she breathed the thin air, then wrapped her arms around her slime-covered unclothed blood dripping body. Looks of astonishment and fear mixed with guilt escaped the space suits of the three, as the single word “help” escaped her lips.

“Abe here. The planet’s breaking apart and its shield is gone. We have a fix on you. Permission to beam you out of there?

Instead of hearing the expected “It’s already been given”, Abe received a nervous reply, “Five to beam back—we’ll help you to get a lock on the human near us. Have our best medic there!” In under a minute, Bill, Thom, Nathaniel, Dennis, and the suffering woman appeared in the transporter room. The explosion on Planet Roxy below grew exponentially in strength.

Kelly Wong

The ship leaped into the Tai stargate, barely escaping the shockwaves set off by Planet Roxy, as she exploded. Nathaniel was right, Roxy was history.

Bill and the others peered out from behind a circular force barrier, which separated the transporter from the rest of the transport room. Bill saw his chief medical officer, Dr. Chelly Manysuns, two of her assistants, and several armed troopers, along with the mysterious beaming officer. He looked around in the transporter, and time slowed as memories of what had just happened flashed before him. He forced himself into the present, where he saw Nathaniel nursing his scorched arm. Next to Nathaniel stood Dennis, his eyes still burning with the recent intensity of battle. Bill’s gaze next fell upon Thom, whose arms were increasingly stained by blood as he supported a miserable form that fought for her every breath. The woman! Bill sprung into action: “Steralization should be complete. Lower the transporter shielding. Thom, she needs the doctor!”

Dr. Manysuns and her two assistants carefully hurried away with their critical patient. The troopers relaxed, and the four in space suits began undressing. When the suits were removed and put away, Thom said wearily, “We all desparately need rest. I recommend we lay in a course for Jash Stargate, where we can gather further intelligence about the extent of the droid invasion. In the meantime, we can get some rest and perhaps learn something from the woman who returned with us.” Everyone concurred. Nathaniel reported to the medic room to have his left arm treated, while Dennis, Thom, and Bill returned to their rooms for some much-needed sleep.

Sleep arrived easily, as did the nightmares. Dennis found himself in the recent battle, but this time each droid he shot suddenly looked just like Dennis and died. Thom wept in his sleep, thinking the woman he rescued had died. In his slumber, Bill sometimes thought that the woman they had found on Roxy below was his lover Tasha. Nathaniel, who was under the influence of painkillers, slept like a digital wristwatch with no batteries.

Exactly six hours after returning from the planet, Bill and the others awoke almost simultaneously to their alarms. Bill eye’s opened and peered out his window, which revealed black streaked stargate-distorted space; he felt the cold window as it dimly hummed. He quickly washed, dressed, and proceeded to the hospital.

“Hello Bill” said Thom with a smile as he saw Bill step into the hospital room.

“You look good” replied Bill. “Any news about our patient?”

Dennis and Nathaniel stepped into the room as Thom answered. “She has made a remarkable recovery. According to Dr. Chelly, the patient’s name is Kelly Wong, and she is human. Beyond that, we know very little.”

Dr. Chelly majestically glided her inhuman form across the room as she added, “Kelly Wong is likely to fully recover her physical health. I didn’t learn much more about her identity before we let her fall into a comfortable sleep. Her injuries and condition suggest that she was connected to some sort of machine and that she had painfully ripped herself out of it during your battle below.”

“How long until she can be safely awakened?” inquired Bill as he cast a glance on the brave slumbering figure of Kelly Wong. “It is of utmost importance that we learn what she knows!”

“Give her at least five more hours,” responded Dr. Chelly. In fact, the doctor recognized her patient, who she had met under dire circumstances several years ago, but she decided it would be best for the health of the crew if she not reveal this information just yet. She wanted some extra time to speak with the patient alone.

Bill agreed. “Alright. Thom, please stay with her. If she wakes up, let me know immediately. Also make sure she doesn’t leave. Her long fight for escape at all costs might not easily leave her mind.” At this news, Thom smiled but Dr. Chelly’s long pointed ears tensed a little.

Bill, Nathaniel, and Dennis proceeded to the bridge. Once they found that everything was in order, they called a meeting in the strategy room. Jerrisons, Nathaniel, Bill, Dennis, and a few other officers arrived. As Bill began summarizing the recent battle, he receives an urgent call from Dr. Chelly. “Excuse me, I’m receiving a call from the doctor.”

The doctor’s distressed voice filled the hall: “Bill, Kelly Wong has vanished!”

Thom’s voice added, ”I haven’t even been able to locate her using the ship’s sensors, and the camera logs seem to have been modified. She was here one moment, I blinked, and she was gone the next.”

Bill’s face contorted in frustration as a word invented by Homer Simpson peirced his lips: “Doh! This is just what I was afraid of. I can’t believe this. Argh!” As frustration transformed itself into arousal, he announced “The meeting is over. Nathaniel and Dennis come with me. It is absolutely imperative that we find Kelly Wong.”

Bill, Nathaniel, and Dennis entered the room, which was recently occupied by Kelly. Thom glanced up for a moment, then turned back to his scans, still baffled by how an entire human being could vanish from his sensors, and on his own ship nonetheless! Maybe, he tried not to think, he also missed her a little bit.

Meanwhile, Kelly Wong relaxed in front of a holographic computer terminal in a forgotten storage area, learning about the Stein ship. No scanners detected her, because she told them not to. The ship’s scanners were old friends. She helped design them while a security apprentice in Hong Kong. Putting backdoors in security systems is irresponsible; so are some people.

Kelly discovered that the Stein ship was purchased by Dennis Stein three years ago using funds from the Stein Foundation. While at the Star Officer academy, Dennis made investments as “hobby”, and was so successful that he was able to buy one of Earth’s top space ships. When Dennis and his class graduated, several formed a private exploration crew on the Stein ship, instead of following the usual route of joining the Space Corps. Privately funded space ships were quite rare, but these independent ships had made some of the most important discoveries of the last few decades. Kelly thought, “Perhaps this is one reason this ship wasn’t captured by the invading droid armies along with Earth’s Space Corps ships.”

Kelly convulsed as memories of the droid armies flooded her mind; her recent injuries ached, and she bled a little. She thougt back on what had happened long ago. Totally by surprise, huge numbers of droids had descended on Hong Kong. Massive droid ships hovered over Victoria Harbor. She remembered those last moments. Guns were discharging everywhere in a Tsim Tsha Tchoi (spelling?) high rise as she battled the Diad Renegade Forces. The battle was the culmination of a carefully orchestrated covert operation. Her right foot cracked a nose as her left hand flew around and crushed a kidney. Then, from nowhere, massive numbers of droids cast a shadow over the city.

Though Kelly destroyed many droids, eventually she fell unconscious. When she awoke, she felt like she was back in Hong Kong, and the droids were gone. In fact, she was lying on a table a vast distance from Hong Kong, giving unwittingly helping the droids! Eight hours ago she figured this out.

Eight hours ago Kelly Wong was encased on a table with hundreds of small wires connected to her body. When Bill and his crew landed on Planet Roxy, Kelly was being fed signals that convinced her that the Diad were attacking her headquarters in Hong Kong. When Bill, Thom, and Nathaniel entered the control area, Kelly was convinced the Diad had entered an important zone, and, in her dellusions, she did everything she could to fight them off. Her combat intelligence was channelled into the droids. Kelly had convulsed a little, while she lay on that table; in her mind she was flipping through the air, kicking, and punching. She had thought she saw her enemy and fought to destroy!

While she was fighting, she thought, one of the droid wires must have snapped. When it had snapped, she conjectured, she must have convulsed and jerked her body even more. This must have snapped more wires until finally she awoke. She thought back to when she had opened her eyes, lifted her head a little off the table, and saw small golden threads protruding from all over her body. She remembed how she felt suffocated by the transparent cover over the table on which she lay, and how, in the distance, she saw men in space suits. She also remembed seeing other tables with people on them like her, some of them part of her team. There was Chan Chai, who had often made her laugh over tea before a dangerous mission. And near him lay John Kurst in droid wires, a huge tough Australian, who always understand her.

When she saw her team members trapped, she knew what to do. She had jerked and ripped the wires away, taken a huge breath, then smashed her arms at the cover. But it didn’t break. She kicked and punched it, but still it didn’t break. Finally, in desparation, she had wrapped the wires around her hand and struck out, smashing her prison wall. She couldn’t remember what happened next, except that she awoke in a space ship.

Kelly Wong smiled, her hands danced around her holographic keyboard. It’s time to rescue my team, she thought with a smile.

Kelly turned again to the computer and searched for information about the droids. She learned that the Stein ship had been captured, but then escaped, and she read about Jerrison’s forces. Then she learned about the mission to destroy the droid manufacturing facility. It had been a success! A success? “No, that can’t be,” she said. A minute later her fist struck the computer hologram terminal, causing her knuckles to bleed. “No way.” Great pain and anger caught fire deep inside her. “There is no way. This is not possible.” For a few minutes she believed that her team was alive, still imprisoned in the droid manufacture facility. She smashed her other hand into the terminal, hard, then curled up and cried like she hadn’t since when her boyfriend had killed himself 10 years ago.

“I found her,” announced Thom. “She’s in cargo bay six.” Thom moved his hand and a holographic display of cargo bay six appeared. Kelly lay curled up in a distant corner. The display zoomed in on her.

“Let’s go!” shouted Bill, “before it’s too late. Come with us Doc.”

Moments later Bill, Thom, Dr. Shelly, and Dennis carefully entered the cargo bay. Bill’s eyes scanned the smooth distant ceiling, then fell on the pitiful girl, who now lay gasping on the floor. He carefully approached her and said, “Kelly Wong, I’m Bill Stein, Captain of the Stein ship. We rescued you from a droid manufacturing facility.”

Slowly, Kelly became aware that she was not alone. Bill’s voice sounded synthetic to her and she felt like droids were everywhere. Weakened, nothing but her instincts supported her, so she fought. She sprung from the floor and flew into the air spinning, her green paper gown barely remaining with her. Her right heal extended and caught Bill squarely in the head, knocking him senseless to the floor. As she landed, with lightning speed she surveyed the field, and saw three figures, which she considered her enemies. One of them, Dennis, raised some kind of stun weapon and fired at her, but she dodged it by flipping backward into the air and landing on a 3 meter black cube that sat behind her.

Thom yelled, “Dennis, look what she did to Bill. We have to neutralize her! She’s crazy!”

Thom and Dennis spread out, as they heard Dr. Shelly shout at Kelly, “Agent Wong, I’m Dr. Shelly Manysuns! We were together during your Mars assignment! You saved my life!!” Thom and Dennis looked at each in surprise, then simultaneously they shot stun rays at Kelly, who easily dodged them as she flipped through the air off of the cube and landed a few meters from the doctor. Kelly stared directly at the doctor, and a glimmer of recognition appeared in her eyes, then Kelly collapsed on the floor, her hands still bleading.

Dennis shot an energy net over Kelly, and Thom approached slowly. The doctor helped Bill, who was dizzy, confused, and had quite a headache. Kelly was carried directly to the doctor’s office, where she was gently set on a table, which was surrounded by a force shield and carefully guarded.

The Stein ship hid in hyperwarp for several days, during which Kelly’s physical wounds healed, and, with Thom’s help, her mind began to come to grips with the horrors she had endured. Meanwhile Bill, Thom, Jerrisons and the others searched for a way to fight the droid invaders.

Droid Space

Fully two weeks had passed while Bill’s crew and Jerrison’s ships lay hidden in hyperwarp. Kelly, dawning a black and silver combat officer’s uniform, took her place as a member of the crew. Thom spent long hours knocking his head against many walls while he searched for a way to successfully attack the droids. Dennis frequently sat on the bridge with a holographic array of weapons spread before him; he searched his mind for a way to address the droid situation. Occassionally, in frustration, he fired off a laser burst into the void. This was safe; Dennis knew what he was doing. Jerrisons kept his troops strong. The rest of the crew went about its business, though the monotony of waiting to do something about the droids slowly extracted a toll. Sometimes they very quietly suggested that their leadership wasn’t up to the challenge. Though heavily armed, the Stein ship was primarily an exploration vehicle, not a battle cruiser, and the crew were correspondingly accustomed to frequent low level excitement rather than long waits interrupted by brief life or death situations.

Out of these long frustrating days and the input of many ship members, a plan slowly formed. When the plan crystallized, Bill gathered the officers in the medieval strategy room. He leaned against a wall of cold granite, which had been excavated from the mountains of northwestern America, and looked around at his officers, his ally Jerrisons, and the dangerous Officer Wong.

Bill stepped forward and said, “When we were detained on Earth, we escaped because Thom was able to control, to some extent, a droid that guarded us.” He turned towards Thom, then looked around the room, and continued, “Thom believes that there is a vast network that links the droids together. This must be how they communicate and distribute the intelligence they parasitically draw from their human and other hosts. This network, which Thom calls Droid Space, is the droid consciousness, the droid reality. Thom and I hope to gain access to Droid Space, search its vast recesses for what is guiding the droids, and reprogram it, so that the droids can be defeated.”

Bill shot a penetrating glance at Dennis, then Kelly, and finally Jerrisons. “To access Droid Space, we’ll have to leave hyperwarp and fight our way into heavily occupied droid territory. If Kelly’s experience is any guide, the droids will be smarter now than ever before, as they have had more time to feed on their captives’ knowledge. Thom and I will enter a capsule like Office Wong was imprisoned in, and, using special equipment developed by the doctor, we will enter their space and fight back. Meanwhile, you must keep us safe. You will receive your orders shortly.” The meeting ended; the final battle had just begun.

Thom’s thoughts represented those of the crew, as he shivered with enthusiasm about the impending battle. A mix of emptiness and anger had grew inside him, feeding on the recent loss. Nothing had been done, no real attack had been launched, and his friends and relatives were likely “food for droid thought”, just as Officer Wong had been! Why had the droids attacked? What had anyone done to provoke them? And as they waited and planned, the droids only grew stronger as they methodically crushed millions of years of progress. He his stung as he smashed his holographic controls. The controls immediately sprang back into existence. Could he?

* * *

A woman carefully placed a silvery leaf beneath a chromatic circular instrument. Satisfaction graced her face as a three-dimensional rotating image of a tiny speck on the leaf appeared above her instrument, magnified 100000 times. She grined as she imagined that the tiny organism smiled as it is was bathed in scanning particles.

Hours later Manir, chief of security, nodded his head at her. “Tasha, two years of hard search, and finally you’ve found your answer in one of Dinesta’s remotest rain forests.”

The Universe is a dangerous place. Scientists don’t travel into the interstellar field without a security force, though often the scientists view security as nothing but a nuisance. Not Tasha. She trusted Manir deeply; he was of her tribe. Those who long ago thought human tribes would vanish, as Earth’s races melted in a pot, were wrong. As old grudges vanished, the inspirational and focusing power of tribal chants returned, and certain tribes grew to dominate planetary science, or planetary medicine as they liked to call it.

“Perhaps now we can return to our Earth? I am deeply worried. We have received no transmissions from there recently...”

“And your incessant attempts to contact Earth have failed misserably,” quipped Tasha, “as they must. The Dinestian Chriono-ellipsoid interrupts interstellar communication.”

Manir cocked his head as he received a signal from his nearby defense monitoring station. “Something is approaching.”

* * *

Bus, the Supreme Droid Commander, strolled with vibrating confidence towards the view portal in his ship. Hundreds of years of hard work had finally paid dividends. A smile broke into his face, as he surveyed the huge planet below. He congratulated himself as he gazed upon thousands of selfreplicating droid-manufacture stations, which visibly spread across the huge planet Peptune, feeding on the massive energy supplies hidden under the surface. He cringed as he recalled the catostrophic failure that had destroyed Roxy, plainly illustrating the danger of an overly connected power grid. “Fortunately that explosion destroyed the Stein ship,” he heard himself say, but then added “the absence of proof is not proof of absence.”

A droid, this one so human looking that it could be mistaken for one, strolled into his quarters, and said quite naturally “It is time.” Bus followed the droid into a hall way, down an elevator (down being defined by the artifical gravity), and through another long hallway. As he walked Bus let his mind fill with difficult millitary problems that his droid army currently faced. The Hartzones and Zartans were currently mounting a combined resistance on a part of Peptune that hadn’t yet been “liberated”. Intelligence suggested that there might be some powerful uncaptured Earth ships scattered through nearby galaxies, like a few needles in an ocean, waiting to awaken and painfully prick. With anticipation, Bus entered a large room.

Before him were three tables, and atop each table a small clear chamber like the one that had imprisoned Kelly Wong. The first chamber held a Hartzone leader, the second a prominent Zartan general, and the third a famous human applied mathematician, each covered like porcupines in wires. Bus set a small device on his head, and the Hartzone leader suddenly awakened and believed himself to be at vigorous war with the Zartans. Terrified of the prospects of a horrible defeat, the Hartzone leader found a way out; he didn’t hate the Zartans, but he loved the Hartzones, and couldn’t bare the possibility of their destruction. So he made a precise order to exploit a Zartan weakness, recently discoverd by his intelligence forces, which could be used to utterly destory the Zartans. And as the order was made, Bus listened to every detail.

Next, Bus awoke the Zartan general, and quickly learned how to completely destroy the Hartzones. Finally, his attention came to rest on the Earth mathematician. A flick of a finger, and the mathematician awoke, and was plagued with the strongest curiosity to know exactly where Earth’s scientific exploration missions were. But wait, he mumbled to himself, why would I, Sol Levinson, care about some nutty bioligists out in the field? Maybe there is a pattern in their locations, thought Sol. He looked around, and found himself on board a ship, wondering how to find the locations of the scientific missions, then wondering why he was on a ship (he hated to fly), and still wondered why he would care about science field trips, and what went on in those missions anyways. This reminded him of some work of Tasha Bowarm (she’s on a field trip to Dinestia...) on fluid flows at the subatomic level that he had begun to generalize. His ideas began flooding his mind, and suddenly all went dim, as Bus ripped off his helm and fainted to the floor.

After regaining his strength, which felt like recovering from the worst possible nightmare, Bus recalled all that he had learned, and shot back to his observation and command deck full of strength and purpose. He knew what had to be done.

* * *